OpenStudy (anonymous):

the area of a square garden is 200m^2. how long is the diagonal?

OpenStudy (anonymous):

help help help help!!!!!!!!!! i need help...ugh!

OpenStudy (anonymous):

area of a square = L*L 200 = L*L l = 14.14

OpenStudy (anonymous):

so diagonal^2 = l^2 + l^2 d^2 = 200 +200 d^2 = 400 d = 20